Find the time required for a 50 Hz alternating current to change its value form zero to the rms value.

Grade:10

1 Answers

Arun Kumar IIT Delhi
10 years ago
Hello Student,
\\v(t)=vsin(2\pi 50t) \\v_{rms}={ v \over \sqrt2} \\=>$so v(t)=0 at t=0$ \\=>v(t)={v \over \sqrt2} \\=>2\pi 50t={\pi \over 4} \\=>t={1 \over 100}
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty